I need help finding the solution to this problem.
I know that the answer is 5, but I'm not sure how to get it.
> 10p+6=12p-4

Answer by gina(6)

10p + 6 = 12p - 4
10p - 10p + 6 = 12p - 10p -4
6 = 2p - 4
6 + 4 = 2p - 4 + 4
10 = 2p
10/2 = 2p/2
5 = p